# The Computational Complexity of Circuit Discovery for Inner Interpretability

### v2

@@ Abstract
Many proposed applications of neural networks in machine learning, cognitive/brain science, and society hinge on the feasibility of inner interpretability via circuit discovery. This calls for empirical and theoretical explorations of viable algorithmic options. Despite advances in the design and testing of heuristics, there are concerns about their scalability and faithfulness at a time when we lack understanding of the complexity properties of the problems they are deployed to solve. To address this, we study circuit discovery with classical and parameterized computational complexity theory: (1) we describe a conceptual scaffolding to reason about circuit finding queries in terms of affordances for description, explanation, prediction and control; (2) we formalize a comprehensive set of queries that capture mechanistic explanation, and propose a formal framework for their analysis; (3) we use it to settle the complexity of many query variants and relaxations of practical interest on multi-layer perceptrons (part of, e.g., transformers). Our findings reveal a challenging complexity landscape. Many queries are intractable (NP-hard, $\Sigma^{p}_{2}$-hard), remain fixed-parameter intractable (W[1]-hard) when constraining model/circuit features (e.g., depth), and are inapproximable under additive, multiplicative, and probabilistic approximation schemes. To navigate this landscape, we prove there exist transformations to tackle some of these hard problems (NP- vs. $\Sigma^{p}_{2}$-complete) with better-understood heuristics, and prove the tractability (PTIME) or fixed-parameter tractability (FPT) of more modest queries which retain useful affordances. This framework allows us to understand the scope and limits of interpretability queries, explore viable options, and compare their resource demands among existing and future architectures.
@@ 1 Introduction
Circuit discovery and its challenges. Since top-down approaches to inner interpretability (see Vilas et al., 2024a) work their way down from high-level concepts or algorithmic hypotheses (Lieberum et al., 2023), there is interest in a complementary bottom-up methodology: circuit discovery (see Shi et al., 2024; Tigges et al., 2024). It centers around neuron- and circuit-level isolation or description (e.g., Hoang-Xuan et al., 2024; Lepori et al., 2023) and attempts to build up higher-level abstractions from this low-level foundation. The motivation is the circuit hypothesis: that models implement their capabilities via small subnetworks (Shi et al., 2024). Advances in the design and testing of interpretability heuristics (see Shi et al., 2024; Tigges et al., 2024) come alongside interest in the automation of circuit discovery (e.g., Conmy et al., 2023; Ferrando & Voita, 2024; Syed et al., 2023) and at the same time concerns about its feasibility (Voss et al., 2021; Räuker et al., 2023). One issue is the challenge of scaling up methods to larger networks, more naturalistic datasets, and more complex tasks (e.g., Lieberum et al., 2023; Marks et al., 2024), given their search over large search spaces involving some manual-intensive steps (Voss et al., 2021). A related issue is that current heuristics, though sometimes promising (e.g., Merullo et al., 2024), often yield discrepant results (see e.g., Shi et al., 2024; Niu et al., 2023; Zhang & Nanda, 2023). They often find circuits that are not functionally faithful (Yu et al., 2024a), or that lack the expected affordances (e.g., effects on behavior; Shi et al., 2024). This questions whether certain localization methods yield results that inform editing (Hase et al., 2023), and vice versa (Wang & Veitch, 2024). More broadly, we run into ‘interpretability illusions’ (Friedman et al., 2024) when our simplifications (e.g., circuits) mimic the local input-output behavior of the system but lack global faithfulness (Jacovi & Goldberg, 2020) .
@@
Exploring viable algorithmic options. These challenges come at a time when, despite emerging theoretical frameworks (e.g., Vilas et al., 2024a; Geiger et al., 2024), there are notable gaps in the formalization and analysis of the computational problems that interpretability heuristics attempt to solve (see e.g., Wang & Veitch, 2024, §8). Issues around scalability of circuit discovery and faithfulness have a natural formulation in the language of Computational Complexity Theory (Arora & Barak, 2009; Downey & Fellows, 2013). A fundamental source of breakdown of scalability — which lack of faithfulness is one manifestation of — is the intrinsic resource demands of interpretability problems. In order to design efficient and effective solutions, we need to understand the complexity properties of circuit discovery queries and the constraints that might be leveraged to yield the desired results. Though a decade of experimental efforts has made promising inroads, the complexity-theoretic properties that naturally impact scalability and faithfulness remain open questions (see e.g., Subercaseaux, 2020, §6C). We settle them here by complementing these efforts with a systematic study of the computational complexity of circuit discovery for inner interpretability. We present a framework that allows us to (a) understand the scope and limits of interpretability queries for description/explanation and prediction/control, (b) explore viable options, and (c) compare their resource demands among existing and future architectures.
@@ 1.2 Overview of results
  • We prove there exist transformations to tackle some of these hard problems (NP- vs. $\Sigma^{p}_{2}$-complete) with better-understood heuristics, and prove the tractability (PTIME) or fixed-parameter tractability (FPT) of other queries of interest, and identify open problems.
  • We describe a quasi-minimality property of ANN circuits and exploit it to generate tractable queries which retain useful affordances, and efficient algorithms to compute them.
  • We establish a separation between local and global query complexity. Together with quasi-minimality, they explain interpretability illusions of faithfulness observed in experiments.
@@ 1.3 Related work
This paper gives the first systematic exploration of the computational complexity of inner interpretability problems. An adjacent area is the complexity analysis of explainability problems (Bassan & Katz, 2023; Ordyniak et al., 2023). It differs from our work in its focus on input queries — aspects of the input that explain model decisions — whereas we look at the inner workings of neural networks via circuit queries. Barceló et al. (2020) study the explainability of multi-layer perceptrons compared to simpler models through a set of input queries. Bassan et al. (2024) extend this idea with a comparison between local and global explainability. None of these works formalize or analyze circuit queries (although Subercaseaux, 2020, identifies it as an open problem); we adapt the local versus global distinction in our framework and show how our proof techniques can tighten some results on explainability queries. Ramaswamy (2019) explores a small set of circuit queries and only on abstract biological networks modeled as general graphs, which cannot inform circuit discovery in ANNs. More generally, we join efforts to build a solid theoretical foundation for interpretability (Bassan & Katz, 2023; Geiger et al., 2024; Vilas et al., 2024a).
@@ 2 Mechanistic understanding of neural networks
Mechanistic understanding is a contentious topic (Ross & Bassett, 2024), but for our purposes it will suffice to adopt a pragmatic perspective. In many cases of practical interest, we want our interpretability methods to output objects that allow us to, in some limited sense, (1) describe or explain succinctly, and (2) control or predict precisely. Such objects (e.g., circuits) should be ‘efficiently queriable’; they are often referred to as “a way of making an explanation tractable” (Cao & Yamins, 2023). Roughly, this means that we would like short descriptions (e.g., small circuits) with useful affordances (e.g., to readily answer questions and perform interventions of interest). Circuits have the potential to fulfill these criteria (Olah et al., 2020). Here we preview some special circuits with useful properties which we formalize and analyze later on. Table 1 maps the main circuits we study to their corresponding affordances for description, explanation, prediction and control. Formal definitions of circuit queries are given alongside results in Section 4 (see also The Computational Complexity of Circuit Discovery for Inner Interpretability).
@@
Sufficient Circuit Which neurons suffice in isolation to cause a behavior? Minimum: shortest description possible in the model. Inference in isolation. Minimal: ablating any neuron breaks behavior of the circuit.
@@
Circuit Robustness How much redundancy supports a behavior? How resilient is it to perturbations? Ablating any set of neurons of size below threshold does not break behavior.
Patched Circuit Which neurons drive a behavior in a given input context? Control nodes? Patching neurons changes network behavior for inputs of interest. Steering; Editing.
@@ 3 Inner interpretability queries as computational problems
3.1 Query variants: coverage, size and minimality
The coverage of a circuit refers to the domain over which it behaves in a certain way (e.g., faithful to the model’s prediction). Local circuits do so over a restricted set of inputs. Global circuits do so over all possible inputs. The size of a circuit is measured in number of neurons. Some circuit queries will require circuits of bounded size whereas others leave the size unbounded. A circuit with a certain property (e.g., local sufficiency) is minimal if there is no subset of its neurons that also has that property (not to be confused with minimum size among all such circuits present in the network). To fit our comprehensive suite of problems, we explain how to generate problem variants and later on only present one representative definition of each. (We simplify without loss of generality).
@@ 3.1 Query variants: coverage, size and minimality
    a Property circuit $\mathcal{C}$ in $\mathcal{M}$ of SizeOUT, such that
    CoverageOUT $\mathcal{C}(\mathbf{x})=\mathcal{M}(\mathbf{x})$, Suffix.
Input: and Table 2 illustrate how to generate problem variants using a template, and ProblemName = Sufficient Circuit as an example. Problem definitions will be given for search (return specified circuits) or decision (answer yes/no circuit queries) versions. Others, including optimization (return maximum/minimum-size circuits), can be generated by assigning variables. Problems presented later on are obtained similarly. We also explore various parameterizations, approximation schemes, and relaxations that we explain in the following sections as needed.
@@
CoverageIN an input $\mathbf{x}$ an input $\mathbf{x}$ an input $\mathbf{x}$ “ ” “ ” “ ”
CoverageOUT “ ” “ ” “ ” $\forall_{\mathbf{x}}$ $\forall_{\mathbf{x}}$ $\forall_{\mathbf{x}}$
SizeIN int. $u\leq|\mathcal{M}|$ “ ” “ ” int. $u\leq|\mathcal{M}|$ “ ” “ ”
SizeOUT size $|\mathcal{C}|\leq u$ “ ” min. size size $|\mathcal{C}|\leq u$ “ ” min. size
Property minimal / “ ” minimal / “ ” “ ” minimal / “ ” minimal / “ ” “ ”
Suffix if it exists, otherwise $\bot$ “ ” “ ” if it exists, otherwise $\bot$ “ ” “ ”
3.2 Classical and parameterized complexity
We prove theorems about interpretability queries building on techniques from classical (Garey & Johnson, 1979) and parameterized complexity (Downey & Fellows, 2013). Here we give a brief, informal overview of the main concepts underlying our analyses (see The Computational Complexity of Circuit Discovery for Inner Interpretability for extensive formal definitions). We will explore beyond classical polynomial-time tractability (PTIME) by studying fixed-parameter tractability (FPT). This allows a more fine-grained look at the sources of complexity of problems. NP-hard queries are considered intractable because they cannot be computed by polynomial-time algorithms. However, a relaxation of interest is to allow unreasonable resource demands as long as they are contained in problem parameters that can be kept small in practice. Parameterizing a given neural network and requested circuit leads to parameterized problems (see Table 3 for problem parameters we study later). Parameterized queries in FPT admit fixed-parameter tractable algorithms. W-hard queries, however, do not. We study counting problems via analogous classes #P and #W[1]. We also investigate completeness for NP and classes higher up the polynomial hierarchy such as $\Sigma^{p}_{2}$ and $\Pi^{p}_{2}$ to explore the possibility to tackle hard interpretability problems with better-understood methods for well-known NP-complete problems (de Haan & Szeider, 2017). Most proof techniques involve various kinds of reductions between computational problems.
@@ 3.2 Classical and parameterized complexity
Parameter description Notation
Model (given) Circuit (requested)
@@
Total number of units [^1] Note the colored parameters bound the input size if bounded (hence vacuously making problems tractable). We avoid this in analyses. $\hat{U}=|\mathcal{M}|\leq\hat{L}\cdot\hat{L}_{w}$ $|\mathcal{C}|=\hat{u}$
@@
3.3 Approximation
Although sometimes computing optimal solutions is intractable, it is conceivable we could devise tractable interpretability procedures to obtain approximate solutions that are useful in practice. We consider 5 notions of approximation: additive, multiplicative, and three probabilistic schemes (see The Computational Complexity of Circuit Discovery for Inner Interpretability for formal definitions). Additive approximation algorithms return solutions at most a fixed distance $c$ away from optimal (e.g., from the minimum-sized circuit), ensuring that errors cannot get impractically large ($c$-approximability). Multiplicative approximation returns solutions at most a factor of optimal away. Some hard problems allow for polynomial-time multiplicative approximation schemes (PTAS) where we can get arbitrarily close to optimal solutions as long as we expend increasing compute time (Ausiello et al., 1999). Finally, consider three other types of probabilistic polynomial-time approximability (henceforth 3PA) that may be acceptable in situations where always getting the correct output for an input is not required: (1) algorithms that always run in polynomial time and produce the correct output for a given input in all but a small number of cases (Hemaspaandra & Williams, 2012); (2) algorithms that always run in polynomial time and produce the correct output for a given input with high probability (Motwani & Raghavan, 1995); and (3) algorithms that run in polynomial time with high probability but are always correct (Gill, 1977).
3.4 Model architecture
The Multi-Layer Perceptron (MLP) is a natural first step in our exploration because (a) it is proving useful as a stepping stone in current experimental (e.g., Lampinen et al., 2024) and theoretical work (e.g., Rossem & Saxe, 2024; McInerney & Burke, 2023); (b) it exist as a leading standalone architecture (Yu et al., 2024b), as the central element of all-MLP architectures (Tolstikhin et al., 2021), and as a key component of state-of-the-art models such as transformers (Vaswani et al., 2017); (c) they are of active interest to the interpretability community (e.g., Geva et al., 2022; 2021; Dai et al., 2022; Meng et al., 2024; 2022; Niu et al., 2023; Vilas et al., 2024b; Hanna et al., 2023); and (d) we can relate our findings on the complexity of inner interpretability to those of explainability which also begins with MLPs (e.g., Barceló et al., 2020; Bassan et al., 2024). Although MLP blocks can be taken as a unit to simplify search, it is recommended to investigate MLPs by treating each neuron as a unit (e.g., Gurnee et al., 2023; Cammarata et al., 2020; Olah et al., 2017), as it better reflects the semantics of computations in neural networks (Lieberum et al., 2023, sec. 2.3.1). We adopt this perspective in our analyses. We write $\mathcal{M}$ for an MLP model and $\mathcal{M}(\mathbf{x})$ for its output on input vector $\mathbf{x}$. Its size $|\mathcal{M}|$ is the number of neurons. A circuit $\mathcal{C}$ is a subset of neurons which induce a (possibly end-to-end) subgraph of $\mathcal{M}$ (see The Computational Complexity of Circuit Discovery for Inner Interpretability for formal definitions).
4 Results & Discussion: the complexity of circuit discovery
In this section we present each circuit query with its computational problem and a discussion of the complexity profile we obtain across variants, relaxations, and parameterizations. For an overview of the results for all queries, see Table 4. Proofs of the theorems can be found in the The Computational Complexity of Circuit Discovery for Inner Interpretability.

### v3

@@ Abstract
Many proposed applications of neural networks in machine learning, cognitive/brain science, and society hinge on the feasibility of inner interpretability via circuit discovery. This calls for empirical and theoretical explorations of viable algorithmic options. Despite advances in the design and testing of heuristics, there are concerns about their scalability and faithfulness at a time when we lack understanding of the complexity properties of the problems they are deployed to solve. To address this, we study circuit discovery with classical and parameterized computational complexity theory: (1) we describe a conceptual scaffolding to reason about circuit finding queries in terms of affordances for description, explanation, prediction and control; (2) we formalize a comprehensive set of queries for mechanistic explanation, and propose a formal framework for their analysis; (3) we use it to settle the complexity of many query variants and relaxations of practical interest on multi-layer perceptrons. Our findings reveal a challenging complexity landscape. Many queries are intractable, remain fixed-parameter intractable relative to model/circuit features, and inapproximable under additive, multiplicative, and probabilistic approximation schemes. To navigate this landscape, we prove there exist transformations to tackle some of these hard problems with better-understood heuristics, and prove the tractability or fixed-parameter tractability of more modest queries which retain useful affordances. This framework allows us to understand the scope and limits of interpretability queries, explore viable options, and compare their resource demands on existing and future architectures.
@@ 1 Introduction
Circuit discovery and its challenges. Since top-down approaches to inner interpretability (see Vilas et al., 2024a) work their way down from high-level concepts or algorithmic hypotheses (Lieberum et al., 2023), there is interest in a complementary bottom-up methodology: circuit discovery (see Shi et al., 2024; Tigges et al., 2024). It starts from neuron- and circuit-level isolation or description (e.g., Hoang-Xuan et al., 2024; Lepori et al., 2023) and attempts to build up higher-level abstractions. The motivation is the circuit hypothesis: models might implement their capabilities via small subnetworks (Shi et al., 2024). Advances in the design and testing of interpretability heuristics (see Shi et al., 2024; Tigges et al., 2024) come alongside interest in the automation of circuit discovery (e.g., Conmy et al., 2023; Ferrando & Voita, 2024; Syed et al., 2023) and concerns about its feasibility (Voss et al., 2021; Räuker et al., 2023). One challenge is scaling up methods to larger networks, more naturalistic datasets, and more complex tasks (e.g., Lieberum et al., 2023; Marks et al., 2024), given their manual-intensive search over large spaces (Voss et al., 2021). A related issue is that current heuristics, though sometimes promising (e.g., Merullo et al., 2024), often yield discrepant results (see e.g., Shi et al., 2024; Niu et al., 2023; Zhang & Nanda, 2023). They often find circuits that are not functionally faithful (Yu et al., 2024a) or lack the expected affordances (e.g., effects on behavior; Shi et al., 2024). This questions whether certain localization methods yield results that inform editing (Hase et al., 2023), and vice versa (Wang & Veitch, 2024). More broadly, we run into ‘interpretability illusions’ (Friedman et al., 2024) when our simplifications (e.g., circuits) mimic the local input-output behavior of the system but lack global faithfulness (Jacovi & Goldberg, 2020).
@@
Exploring viable algorithmic options. These challenges come at a time when, despite emerging theoretical frameworks (e.g., Vilas et al., 2024a; Geiger et al., 2024), there are notable gaps in the formalization and analysis of the computational problems that interpretability heuristics attempt to solve (see Wang & Veitch, 2024, §8). Issues around scalability of circuit discovery and faithfulness have a natural formulation in the language of Computational Complexity Theory (Arora & Barak, 2009; Downey & Fellows, 2013). A fundamental source of breakdown of scalability — which lack of faithfulness is one manifestation of — is the intrinsic resource demands of interpretability problems. In order to design efficient and effective solutions, we need to understand the complexity properties of circuit discovery queries and the constraints that might be leveraged to yield the desired results. Although experimental efforts have made promising inroads, the complexity-theoretic properties that naturally impact scalability and faithfulness remain open questions (see e.g., Subercaseaux, 2020, §6C). We settle them here by complementing these efforts with a systematic study of the computational complexity of circuit discovery for inner interpretability. We present a framework that allows us to (a) understand the scope and limits of interpretability queries for description/explanation and prediction/control, (b) explore viable options, and (c) compare their resource demands among existing and future architectures.
@@ 1.2 Overview of results
  • We prove there exist transformations to potentially tackle some hard problems (NP- vs. $\Sigma^{p}_{2}$-complete) with better-understood heuristics, and prove the tractability (PTIME) or fixed-parameter tractability (FPT) of other queries of interest, and we identify open problems.
  • We describe a quasi-minimality property of ANN circuits and exploit it to generate tractable queries which retain useful affordances as well as efficient algorithms to compute them.
  • We establish a separation between local and global query complexity. Together with quasi-minimality, this explains interpretability illusions of faithfulness observed in experiments.
@@ 1.3 Related work
This paper gives the first systematic exploration of the computational complexity of inner interpretability problems.111 This work expands on FA’s PhD dissertation at University of Bristol (Adolfi, 2023; Adolfi et al., 2024). An adjacent area is the complexity analysis of explainability problems (Bassan & Katz, 2023; Ordyniak et al., 2023). It differs from our work in its focus on input queries — aspects of the input that explain model decisions — as we look at the inner workings of neural networks via circuit queries. Barceló et al. (2020) study the explainability of multi-layer perceptrons compared to simpler models through a set of input queries. Bassan et al. (2024) extend this idea with a comparison between local and global explainability. None of these works formalize or analyze circuit queries (although Subercaseaux, 2020, identifies it as an open problem); we adapt the local versus global distinction in our framework and show how our proof techniques can tighten some results on explainability queries. Ramaswamy (2019) and Adolfi & van Rooij (2023) explore a small set of circuit queries and only on abstract biological networks modeled as general graphs, which cannot inform circuit discovery in ANNs. Efforts in characterizing the complexity of learning neural networks (e.g., Song et al., 2017; Chen et al., 2020; Livni et al., 2014) might eventually connect to our work, although a number of differences between the formalizations makes results in one area difficult to predict from those in the other. Likewise, efforts to settle the complexity of finding small circuits consistent with a truth table (Hitchcock & Pavan, 2015) are currently too general to be applicable to interpretability problems. More generally, we join efforts to build a solid theoretical foundation for interpretability (Bassan & Katz, 2023; Geiger et al., 2024; Vilas et al., 2024a).
@@ 2 Mechanistic understanding of neural networks
Mechanistic understanding is a contentious topic (Ross & Bassett, 2024), but for our purposes it will suffice to adopt a pragmatic perspective. In many cases of practical interest, we want our interpretability methods to output objects that allow us to, in some limited sense, (1) describe or explain succinctly, and (2) control or predict precisely. Such objects (e.g., circuits) should be ‘efficiently queriable’; they are often referred to as “a way of making an explanation tractable” (Cao & Yamins, 2023). Roughly, this means that we would like short descriptions (e.g., small circuits) with useful affordances (e.g., to readily answer questions and perform interventions of interest). Circuits have the potential to fulfill these criteria (Olah et al., 2020). Here we preview some special circuits with useful properties which we formalize and analyze later on. Table 1 maps the main circuits we study to their corresponding affordances for description, explanation, prediction and control. Formal definitions of circuit queries are given alongside results in Section 4 (see also Appendix: Definitions, Theorems and Proofs).
@@
Sufficient Circuit Which neurons suffice in isolation to cause a behavior? Minimum: shortest description. Inference in isolation. Minimal: ablating any neuron breaks behavior of the circuit.
@@
Circuit Robustness How much redundancy supports a behavior? Resilience to perturbations. Ablating any set of neurons of size below threshold does not break behavior.
Patched Circuit Which neurons drive a behavior in a given input context, i.e., are control nodes? Patching neurons changes network behavior for inputs of interest. Steering; Editing.
@@ 3 Inner interpretability queries as computational problems
Query variants: coverage, size and minimality. The coverage of a circuit is the domain over which it behaves in a certain way (e.g., faithful to the model’s prediction). Local circuits do so over a finite set of known inputs and global circuits do so over all possible inputs. The size of a circuit is the number of neurons. Some circuit queries require circuits of bounded size whereas others leave the size unbounded. A circuit with a certain property (e.g., local sufficiency) is minimal if there is no subset of its neurons that also has that property (cf. minimum size among all such circuits present in the network; see Figure 1).
Figure 1: Relationships between circuit types. Sufficient Circuits (SCs) are faithful to the model. The entire network is a trivial SC. Necessary Circuits (NCs) are units shared by all minimal SCs. Quasi-minimal SCs contain a known breaking point (here, NC) and unknown superfluous units.
To fit our comprehensive suite of problems, we explain how to generate problem variants and later on only present one representative definition of each.
@@
    A Property circuit $\mathcal{C}$ of $\mathcal{M}$, SizeOUT, s.t. CoverageOUT $\mathcal{C}(\mathbf{x})\!=\!\mathcal{M}(\mathbf{x})$, Suffix.
Input: and Table 2 illustrate how to generate problem variants using a template, and ProblemName = Sufficient Circuit as an example (e.g., the Coverage[IN/OUT] variables specify parts of the input/output description that vary according to whether the requested circuit must have global or local faithfulness). Problem definitions will be given for search (return specified circuits) or decision (answer yes/no circuit queries) versions. Others, including optimization (return maximum/minimum-size circuits), can be generated by assigning variables. Problems presented later on are obtained similarly. We also explore various parameterizations, approximation schemes, and relaxations that we explain in the following sections as needed.
@@
CoverageIN an input $\mathbf{x}$ an input $\mathbf{x}$ an input $\mathbf{x}$ “__” “__” “__”
CoverageOUT “__” “__” “__” $\forall_{\mathbf{x}}$ $\forall_{\mathbf{x}}$ $\forall_{\mathbf{x}}$
SizeIN int. $u\leq|\mathcal{M}|$ “__” “__” int. $u\leq|\mathcal{M}|$ “__” “__”
SizeOUT size $|\mathcal{C}|\leq u$ “__” min. size size $|\mathcal{C}|\leq u$ “__” min. size
Property minimal / “__” minimal / “__” “__” minimal / “__” minimal / “__” “__”
Suffix if it exists, otherwise $\bot$ “__” “__” if it exists, otherwise $\bot$ “__” “__”
3.1 Complexity analyses
Classical and parameterized complexity. We prove theorems about interpretability queries building on techniques from classical (Garey & Johnson, 1979) and parameterized complexity (Downey & Fellows, 2013). Given our limited knowledge of the problem space of interpretability, worst-case analysis is appropriate to explore which problems might be solvable without requiring any additional assumptions (e.g., Bassan et al., 2024; Barceló et al., 2020), and experimental results suggest it captures a lower bound on real-world complexity (e.g., Friedman et al., 2024; Shi et al., 2024; Yu et al., 2024a). Here we give a brief, informal overview of the main concepts underlying our analyses (see Appendix: Definitions, Theorems and Proofs for extensive formal definitions). We will explore beyond classical polynomial-time tractability (PTIME) by studying fixed-parameter tractability (FPT), a more novel and finer-grained look at the sources of complexity of problems to test aspects that possibly make interpretability feasible in practice. NP-hard queries are considered intractable because they cannot be computed by polynomial-time algorithms. A relaxation is to allow unreasonable (e.g., exponential) resource demands to be confined to problem parameters that can be kept small in practice. Parameterizing a given ANN and requested circuit leads to parameterized problems (see Table 3 for problem parameters we study later). Parameterized queries in the class FPT admit fixed-parameter tractable algorithms. W-hard queries (by analogy: to FPT as NP-hard is to PTIME), however, do not. We study counting problems via analogous classes #P and #W[1]. We also investigate completeness for NP and classes higher up the polynomial hierarchy such as $\Sigma^{p}_{2}$ and $\Pi^{p}_{2}$ to identify aspects of hard problems that make them even harder, and to explore the possibility to tackle hard interpretability problems with better-understood methods for well-known NP-complete problems (de Haan & Szeider, 2017). Most proofs involve reductions between computational problems which establish the complexity status of interpretability queries based on the known complexity of canonical problems in other areas.
@@ 3.1 Complexity analyses
Parameter Model (given) Circuit (requested)
@@
Total number of units22footnotemark: 2 $\hat{U}=|\mathcal{M}|\leq\hat{L}\cdot\hat{L}_{w}$ $|\mathcal{C}|=\hat{u}$
@@
[…]
Model architecture. The Multi-Layer Perceptron (MLP) is a natural first step in our exploration because (a) it is proving useful as a stepping stone in current experimental (e.g., Lampinen et al., 2024) and theoretical work (e.g., Rossem & Saxe, 2024; McInerney & Burke, 2023); (b) it exists as a leading standalone architecture (Yu et al., 2024b), as the central element of all-MLP architectures (Tolstikhin et al., 2021), and as a key component of state-of-the-art models such as transformers (Vaswani et al., 2017); (c) it is of active interest to the interpretability community (e.g., Geva et al., 2022; 2021; Dai et al., 2022; Meng et al., 2024; 2022; Niu et al., 2023; Vilas et al., 2024b; Hanna et al., 2023); and (d) we can relate our findings in inner interpretability to those in explainability, which also begins with MLPs (e.g., Barceló et al., 2020; Bassan et al., 2024). Although MLP blocks can be taken as units to simplify search, it is recommended to investigate MLPs by treating each neuron as a unit (e.g., Gurnee et al., 2023; Cammarata et al., 2020; Olah et al., 2017), as it better reflects the semantics of computations in ANNs (Lieberum et al., 2023, sec. 2.3.1). We adopt this perspective in our analyses. We write $\mathcal{M}$ for an MLP model and $\mathcal{M}(\mathbf{x})$ for its output on input vector $\mathbf{x}$. Its size $|\mathcal{M}|$ is the number of neurons. A circuit $\mathcal{C}$ is a subset of $|\mathcal{C}|$ neurons which induce a (possibly end-to-end) subgraph of $\mathcal{M}$ (see Appendix: Definitions, Theorems and Proofs for formal definitions).
4 Results & Discussion: the complexity of circuit queries
In this section we present each circuit query with its computational problem and a discussion of the complexity profile we obtain across variants, relaxations, and parameterizations. For an overview of the results for all queries, see Table 4. Proofs of the theorems can be found in the Appendix: Definitions, Theorems and Proofs.